# Passage-traversing optimal path planning with sampling-based algorithms

### v1

@@ 3.4 Free space decomposition with Gabriel cells
A prominent advantage of the proposed passage detection procedure is that it comes with a natural subdivision of the free space. Delaunay graphs are planar, i.e., edges do not cross each other, and connected without isolated nodes. As a result, the passage distribution essentially defines a decomposition of its enclosed free space. The cells are bounded by passages and have obstacles as their vertices. Together with environment boundaries, a cell decomposition of the entire free space is further obtained. The free space $\mathcal{X}_{free}$ is divided into separate regions $\mathcal{C}_{i}$ such that $\mathcal{X}_{free}=\cup_{i=1}^{n_{c}}\mathcal{C}_{i}$ where $n_{c}$ designates the number of cells. For cell adjacency, $\forall i\neq j$
@@
where $l(\mathbf{p}_{p}^{*},\mathbf{p}_{q}^{*})$ is the neighboring passage segment of $\mathcal{C}_{i}$ and $\mathcal{C}_{j}$. Namely, while obstacle sides comprise parts of cell edges, cells only have passage segments as common edges (see Figure 6). These cells are termed Gabriel cells here. Cells enable many cell-based methods in path planning for coarse solutions (Choset et al., 2005; LaValle, 2006). When integrated into sampling-based PTOPP, they can dramatically speed up passage traversal check of paths via sample localization in cells, making PTOPP computationally efficient without extra decomposition efforts.
Figure 6: Free space is divided into Gabriel cells denoted by circled numbers (environment boundaries are not considered here). Cells are enclosed by passage segments and obstacles. Adjacent cells have passage segments as common edges.
Gabriel cells always contain passage segments as cell sides. Unlike planar graph faces bounded only by graph edges, Gabriel cells additionally have sides from obstacle boundaries in most situations because obstacle sizes are non-negligible as exemplified in Figure 6. Detected passages define a passage graph. To report all cells, face detection is carried out in this graph. Adjacency lists in the passage graph are first sorted counter-clockwise (or clockwise). Let $A_{i}$ be the passage adjacency list of $\mathcal{O}_{i}$, i.e., $A_{i}=\{j\,|\,\mathcal{V}(\mathcal{P}_{i-j})=\text{True}\}$. $A_{i}$ is sorted such that
@@
where vector $\mathbf{c}_{i,j}=\mathbf{c}_{j}-\mathbf{c}_{i}$. Oriented() is a transitivity-preserving comparer telling if it is counter-clockwise to rotate from $\mathbf{c}_{i,j}$ to $\mathbf{c}_{i,k}$. The entire graph is then traversed in a depth-first search (DFS) way in which edges maintain a clockwise ascending order until a face is enclosed. See Algorithm 3 for the Gabriel cell detection procedure. Sorting $O(m)$ items takes $O(m\log m)$ time, and graph traversal takes $O(m)$ time, resulting in overall $O(m\log m)$ complexity. Sequential obstacle indices in cells are recorded in detection. Cell vertices are associated passage segment endpoints between obstacles and are readily retrievable.
Although Gabriel cell decomposition is transplantable to 3D space to decompose the space into voxels in each obstacle height interval, the resulting cell structures have complex boundaries and adjacency relations. To use cells for PTOPP in 3D space, cells on the base are employed. Specifically, planar cells on the base are found considering passages rooted on the base ($\underline{h}=0$). For midair passages with $\underline{h}>0$, their projections on the base intersect with base cells. Since the projection may fall into the obstacle region, obstacles are also treated as cells, and obstacle cross sections are assumed invariant, resulting in the compound Gabriel cell decomposition in Algorithm 4. Finding intersected cells equals the passage traversal check introduced later and takes $O(1)$ time. Midair passage number does not exceed $O(m)$. Therefore, compound cell detection also takes $O(m)$ time. Midair passages are treated as virtual sides of base cells which their projections intersect with. As such, when checking passed passages of paths in 3D space, it suffices to additionally check virtual sides of cells.
1 Get passage adjacency lists $A_{1},...,A_{m}$ from passages;
@@ 14 $u\leftarrow v;e\leftarrow e^{\prime}$;
16return cell list $C$;
@@
1 Detect Gabriel cells using passages starting at base;
@@ 4.1 Minimum passage width PTOPP
which encodes the narrowest passage width passed from $\sigma(0)$ to $\sigma(\tau)$. PTOPP problems solely formulating the minimum passage width in the path cost are classified as minimum passage width PTOPP (MPW-PTOPP). MPW-PTOPP aims to make $f_{p}$ large enough.
Different planning objectives are often conflicting. For instance, a path not traversing narrow paths need not be short and vice versa. Improving one objective usually comes with deteriorating the others (Sakcak and LaValle, 2021). To handle this, a trade-off is often adopted. One versatile cost in (Huang et al., 2024a) is
@@
where $w_{p}>0$ is the weight adjusting the dominance between Len$(\sigma)$ and $f_{p}(\sigma)$, offering (13) good adjustability over other trade-off costs, e.g., the ratio cost $f_{c}(\sigma)=\text{Len}(\sigma)/f_{p}(\sigma)$ in (Huang et al., 2023). Path cost functions need to have certain properties to be usable. A path cost must be monotonically non-decreasing and bounded in path concatenation to be optimizable by sampling-based optimal planners (Karaman and Frazzoli, 2011). To achieve monotonicity of MPW-PTOPP costs, $f_{p}(\sigma)$ is initialized as a large value when $\sigma$ has not passed passages yet, i.e.,
@@
Clearly, $f_{c}(\cdot)$ above satisfies boundedness and monotonicity (Len()($\cdot$)( ⋅ ) is monotonically increasing, and $f_{p}(\cdot)$ is monotonically non-increasing). However, it proves to be insufficient. The cost is still not asymptotically optimizable by sampling-based optimal planners for not being order-preserving. Formally, suppose path $\sigma_{1}$ and $\sigma_{2}$ reach the same endpoint, i.e., $\sigma_{1}(1)=\sigma_{2}(1)$. Both are concatenated with path $\sigma_{3}$ at $\sigma_{3}(0)=\sigma_{1}(1)=\sigma_{2}(1)$. A path cost $f_{c}(\cdot)$ is said to be order-preserving if the following holds
@@
(Pearl, 1984), which can be regarded as a generalized and less demanding condition of the broadly used additivity property of path costs. Sampling-based optimal planners implicitly rely on the order-preserving property of costs to find the optimal solution. A basic hypothesis is that reducing the cost of an intermediate path segment decreases the entire path’s cost. This is utilized in optimal path search in roadmaps, e.g., optimal PRM (PRM), and locally rewiring trees, e.g., RRT. Suboptimal paths are returned for non-order-preserving path costs. To remedy this, a path cost in PTOPP is said to be compatible if it is monotonically non-decreasing, bounded, and order-preserving under path concatenation.
The compound trade-off cost in (13) is not order-preserving due to the interaction between Len$(\sigma)$ and $w_{p}f_{p}(\sigma)$. This can be noticed by investigating the occasion where $f_{p}(\sigma)$ is achieved at $\sigma(\tau_{p})$. To be globally optimal, the optimal path segment from $\mathbf{x}_{0}$ to $\sigma(\tau_{p})$ is the shortest path to attain the minimum Len$(\sigma)$, which, nevertheless, generally cannot be found via optimizing (13). See Figure 7 for an example and conceptual illustration. An order-preserving cost for MPW-PTOPP is directly taking negative $f_{p}(\sigma)$, i.e.,
@@
While $f_{c}(\cdot)$ above is piece-wise constant since it gets updated only if $\sigma$ has passed a passage narrower than before, it is easy to verify that the cost is compatible with sampling-based optimal planners. Paths traversing the same minimum-width passage have an equal cost, leading to a non-zero measure of optimal paths. Let $\Sigma^{*}$ be the set of optimal paths and $\mathcal{X}^{*}$ be the set of states that an optimal path in $\Sigma^{*}$ passes through
@@
and $\mu(\mathcal{X}^{*})$ denotes the hypervolume of $\mathcal{X}^{*}$. A widely used assumption in OPP is zero-measure of optimal paths, i.e., $\mu(\mathcal{X}^{*})=0$, to imply that an optimal path cannot be found in a finite number of iterations. For discretely updated costs (16) in PTOPP, $\mu(\mathcal{X}^{*})>0$ as $\mu(\mathcal{X}^{*})\geq\text{Len}(\sigma^{*})\mu(\mathcal{P}^{*})>0$ where $\mu(\mathcal{P}^{*})>0$ is the hypervolume of the minimum-width passage in dim$(\mathcal{X})$ 1 dimensional space.
Figure 7: Example of lack of optimizability for non-compatible costs by sampling-based optimal planners. Consider cost (13) with a large weight $w_{p}$ and passages $\|\mathcal{P}_{1-2}\|>\|\mathcal{P}_{2-3}\|=\|\mathcal{P}_{4-5}\|$. $\mathbf{x}_{m}$ is an intermediate path node. The globally optimal path is the joint path $\sigma*=\sigma_{2}|\sigma_{3}$, but $\sigma_{1}|\sigma_{3}$ will be found since the optimal path from $\mathbf{x}_{0}$ to $\mathbf{x}_{m}$ is found as $\sigma_{1}$.
A positive $\mu(\mathcal{X}^{*})$ infers a positive probability of finding an optimal path even after a finite number of sample iterations. But returned optimal paths can differ significantly in length and topology properties from run to run by solely maximizing $f_{p}(\sigma)$. To remove this indeterminacy and enhance optimality, extra planning objectives are introduced while the optimal path cost is imposed as a constraint. In particular, for (16), the optimal path is found in a length-regulated form as
$\begin{split}\sigma^{*}&=\operatorname*{arg\,min}_{\sigma\in\mathcal{H}}\,\text{Len}(\sigma)\\ \operatorname{s.t.}\;&f_{c}(\sigma^{*})=\min f_{c}(\sigma)\\ \end{split}$ (18)
where $\mathcal{H}$ is the solution path set for $\min f_{c}(\sigma)$ returned by planners. $\mathcal{X}^{*}$ now has zero measure due to minimization of Len$(\sigma^{*})$. Though $\mu(\mathcal{X}^{*})=0$ does not imply a single optimal path, the stochasticity of the optimal path is removed. $f_{c}(\cdot)$ has a higher priority over Len()($\cdot$)( ⋅ ) and hence the global optimality. The optimal path is reachable as both $f_{c}(\cdot)$ and Len()($\cdot$)( ⋅ ) are compatible. As a result, the compatibility of this cost with sampling-based optimal planners is ensured. In fact, MPW-PTOPP finds paths homotopic to those in maximum clearance path planning because the clearance of a path is half $f_{p}(\sigma)$ in planner cases.
@@ 4.2 Global passage width PTOPP
Despite the simplicity and versatility, only taking into account $f_{p}(\sigma)$ in PTOPP may be insufficient to reach high-quality paths for not considering traversed passages other than $f_{p}(\sigma)$. For example, $f_{p}(\sigma)$ is an inevitable passage from the start position $\mathbf{x}_{0}$ in Figure 8. MPW-PTOPP essentially degrades to shortest path planning. This further scales to occasions where feasible paths have inevitable passages. A typical situation is enclosure of the start by surrounding passages. Feasible paths must traverse one surrounding passage (see Figure 8). Outer passages wider than them are muted in costs, highlighting the limitation of MPW-PTOPP. To alleviate this degradation, a more comprehensive coverage of $P_{\sigma}$ is required. Conceptually, if traversed passage widths more than $f_{p}(\sigma)$ in $P_{\sigma}(1)$ are incorporated in path costs, PTOPP problems are called global passage width PTOPP (GPW-PTOPP).
Figure 8: On the left, the start position is placed on one side of the narrowest passage, and the target is on the other side. Feasible paths must traverse this passage. The start position on the right is enclosed by passages. Any feasible path will pass one of them.
Cost formalisms in GPW-PTOPP are non-trivial due to the involvement of multiple passages. Some common measures are not directly applicable as costs. For instance, maximizing the sum of passed passage widths incurs an ergodic passage traversal, and the average passage width lacks monotonicity. For better comparability of traversed passage widths, denote $\mathbf{p}_{\sigma}(\tau)$ the passed passage widths in ascending order and $\bar{P}_{\sigma}(\tau)$ the sorted permutation of $P_{\sigma}(\tau)$ such that $\|\bar{P}_{\sigma}(\tau,i)\|\leq\|\bar{P}_{\sigma}(\tau,j)\|$ for $i<j$, then
@@
where $\mathbf{p}_{\sigma}(\tau,i)=\|\bar{P}_{\sigma}(\tau,i)\|$. To involve all these passage widths in costs, the path cost summing up all traversed passage widths is framed as
@@
where $\mathbf{w}_{p}\succeq\mathbf{0}$ with nonnegative entries is the weight vector. How $f_{c}(\sigma)$ updates is the key to compatibility, but it is not well-defined in path concatenation since the dimension of $\mathbf{p}_{\sigma}$ will generally increase. An important step is to adopt a fixed dimension of $\mathbf{p}_{\sigma}$ larger than the number of passages a path may pass. Say the dimension is q, $\mathbf{p}_{\sigma}$ is initialized as $\bar{\varepsilon}_{p}\mathbf{1}_{q}$ where $\bar{\varepsilon}_{p}$ is a large value as in (14). $\mathbf{1}_{q}$ is the q-dimensional all-one vector. Every sorted passage width is now assigned a weight. The elements in $\mathbf{w}_{p}$ are in descending order so that narrower passages have higher priorities to be avoided, i.e., $\mathbf{w}_{p}=[w_{p,1},w_{p,2},...,w_{p,q}]^{\mathsf{T}}$, $w_{p,i}<w_{p,i+1}$.
While $\mathbf{p}_{\sigma}$ now preserves its dimensionality and (20) is monotonically non-increasing, it is easy to verify that (20) is not order-preserving as a sum of paired $w_{p,i}\mathbf{p}_{\sigma}(1,i)$. To resolve this, the lexicographic order is utilized for comparing $\mathbf{p}_{\sigma}$ and $f_{c}(\cdot)$ is specially designed to preserve this order via configuring $\mathbf{w}_{p}$. Formally, if $\mathbf{p}_{\sigma_{1}}$ is lexicographically smaller than $\mathbf{p}_{\sigma_{2}}$, then
@@
When $\mathbf{p}_{\sigma_{1}}\prec_{\text{lex}}\mathbf{p}_{\sigma_{2}}$, the narrowest passage $\sigma_{2}$ passes is wider than $\sigma_{1}$ except those common passages for $\sigma_{1},\sigma_{2}$. A smaller cost is thus expected for $f_{c}(\sigma_{2})$. This can be done by setting large differences between successive $\mathbf{w}_{p}$ items so that $w_{p,i}/w_{p,i+1}\gg 1$. For a given passage distribution, it is shown that by setting sufficiently large differences between $w_{p,i}$, the following holds
@@
In practice, q can be fairly small to consider top k narrowest passages and such problems are called k-minimum passage width PTOPP (k-MPW-PTOPP). (20) has a general form of (16) that is able to optimize multiple passed passages globally. Directly composing a compatible trade-off cost analogous to (13) is attainable as
$f_{c}(\sigma)=\text{Len}(\sigma)-\mathbf{w}^{\mathsf{T}}\mathbf{p}_{\sigma}(1).$ (24)
However, to be order-preserving, weights in $\mathbf{w}$ are set equal, making it unable to prioritize avoiding narrower passages (see Appendix A.2 for more analysis of compatible GPW-PTOPP path cost design).

### v2

@@ 3.4 Free space decomposition with Gabriel cells
A prominent advantage of the proposed passage detection procedure is that it brings a natural free space subdivision. Delaunay graphs are planar (i.e., no intersecting edges) and connected (i.e., no isolated nodes). As a result, the passage distribution essentially defines a decomposition of the free space. The cells are bounded by passages and have obstacles as their vertices. Together with environment boundaries, a cell decomposition of the entire free space is obtained. The free space $\mathcal{X}_{free}$ is divided into separate regions $\mathcal{C}_{i}$ such that $\mathcal{X}_{free}=\cup_{i=1}^{n_{c}}\mathcal{C}_{i}$ where $n_{c}$ designates the number of cells. For cell adjacency, $\forall i\neq j$
@@
where $l(\mathbf{p}_{p}^{*},\mathbf{p}_{q}^{*})$ is the neighboring passage segment of $\mathcal{C}_{i}$ and $\mathcal{C}_{j}$. Namely, obstacles and passage segments comprise cell edges, but cells only have passage segments as common edges (see Figure 6). These cells are termed Gabriel cells here, which enable many cell-based path planning methods (choset2005principles; lavalle2006planning). When integrated into sampling-based PTOPP, they can dramatically speed up passage traversal check of paths via sample localization in cells, making PTOPP computationally efficient without extra decomposition steps.
Figure 6: Free space is divided into Gabriel cells denoted by circled numbers (environment boundaries are not considered). Cells are enclosed by passage segments and obstacles. Adjacent cells have passage segments as common edges.
Gabriel cells always contain passage segments as sides. They additionally have sides from obstacle boundaries in most situations for non-negligible obstacle sizes, e.g., Figure 6. Detected passages define a passage graph. To report all cells, face detection is carried out in this graph. Adjacency lists in the passage graph are first sorted counter-clockwise (or clockwise). Let $A_{i}$ be the passage adjacency list of $\mathcal{O}_{i}$, i.e., $A_{i}=\{j\,|\,\mathcal{V}(\mathcal{P}_{i-j})=\text{True}\}$. $A_{i}$ is sorted such that
@@
and $\mathbf{c}_{i,j}=\mathbf{c}_{j}-\mathbf{c}_{i}$. Oriented() is a transitivity-preserving comparer telling if it is counter-clockwise from $\mathbf{c}_{i,j}$ to $\mathbf{c}_{i,k}$. Later the passage graph is traversed in a depth-first search (DFS) way in which edges maintain a clockwise order until a face is enclosed. Cell vertices are associated passage segment endpoints and are readily retrievable. See Algorithm 3 for the Gabriel cell detection procedure. Sorting $O(m)$ items takes $O(m\log m)$ time, and graph traversal takes $O(m)$ time, resulting in overall $O(m\log m)$ complexity.
Although Gabriel cell decomposition is transplantable to 3D space to decompose the space into voxels in obstacle height intervals, the resulting cells have complex boundaries and adjacency relations. To use cells for PTOPP in 3D space, cells on the base are employed. Specifically, planar cells on the base are found considering passages rooted on the base ($\underline{h}=0$). For midair passages with $\underline{h}>0$, their projections on the base intersect with base cells. Since the projection may fall into the obstacle region, obstacle region is treated as planar cells (obstacle cross sections are assumed invariant), resulting in compound Gabriel cell decomposition in Algorithm 4. Finding intersected cells takes $O(1)$ time by calling the passage traversal check routine. Midair passage number does not exceed $O(m)$. Therefore, compound cell detection also takes $O(m)$ time. Midair passages are treated as virtual sides of base cells which their projections intersect with. When checking passed passages of paths in 3D space, virtual sides of cells are additionally checked.
1Get passage adjacency lists $A_{1},...,A_{m}$ from passages;
@@ 13 $u\leftarrow v;e\leftarrow e^{\prime}$;
15return cell list $C$;
@@
1Detect Gabriel cells using passages starting at base;
@@ 4.1 Minimum passage width PTOPP
which encodes the narrowest passage width passed from $\sigma(0)$ to $\sigma(\tau)$. Problems solely formulating the minimum passage width in the path cost are classified as minimum passage width PTOPP (MPW-PTOPP). MPW-PTOPP aims to make $f_{p}$ large enough.
Different planning objectives are often conflicting. For instance, a path not traversing narrow paths need not be short and vice versa. Improving one objective usually comes with deteriorating the others (sakcak2021complete). To handle this, a trade-off is often adopted. One versatile cost used in (huang2024homotopic) is
@@
[…]
@@
Clearly, $f_{c}(\cdot)$ above satisfies monotonicity and boundedness (Len()($\cdot$) is monotonically increasing. $f_{p}(\cdot)$ is monotonically non-increasing). However, it proves to be insufficient. The cost is still not asymptotically optimizable by sampling-based optimal planners for not being order-preserving. Formally, suppose path $\sigma_{1}$ and $\sigma_{2}$ reach the same endpoint, i.e., $\sigma_{1}(1)=\sigma_{2}(1)$. Both are concatenated with path $\sigma_{3}$ at $\sigma_{3}(0)=\sigma_{1}(1)=\sigma_{2}(1)$. A path cost $f_{c}(\cdot)$ is said to be order-preserving if the following holds
@@
(pearl1984heuristics), which can be regarded as a generalized and less demanding condition of the broadly used additivity property of path costs. Sampling-based optimal planners implicitly rely on the order-preserving property of costs to find the optimal solution. A basic hypothesis is that reducing the cost of an intermediate path segment decreases the entire path’s cost. This is utilized in optimal path search in roadmaps, e.g., optimal PRM (PRM), and locally rewiring trees, e.g., RRT. Suboptimal paths are returned for non-order-preserving path costs. To remedy this, a path cost in PTOPP is said to be compatible if it is monotonically non-decreasing, bounded, and order-preserving under path concatenation.
[…]
@@
While $f_{c}(\cdot)$ above is piece-wise constant since it gets updated only if $\sigma$ has passed a passage narrower than before, it is easy to verify that the cost is compatible. Paths traversing the same minimum-width passage have an equal cost, leading to a non-zero measure of optimal paths. Let $\Sigma^{*}$ be the set of optimal paths and $\mathcal{X}^{*}$ be the set of states that an optimal path in $\Sigma^{*}$ passes through
@@
[…]
Figure 7: Example of lack of optimizability for non-compatible costs by sampling-based optimal planners. Consider cost (14) with a large weight $w_{p}$ so $f_{p}(\sigma)$ dominates the cost. Passages $\|\mathcal{P}_{1-2}\|>\|\mathcal{P}_{2-3}\|=\|\mathcal{P}_{4-5}\|$. $\mathbf{x}_{m}$ is an intermediate path node. The globally optimal path is the joint path $\sigma*=\sigma_{2}|\sigma_{3}$, but $\sigma_{1}|\sigma_{3}$ is found since the optimal path from $\mathbf{x}_{0}$ to $\mathbf{x}_{m}$ is found as $\sigma_{1}$.
A positive $\mu(\mathcal{X}^{*})$ infers a positive probability of finding an optimal path even after a finite number of sampling iterations. But returned optimal paths differ significantly in terms of length and topology properties from run to run by solely maximizing $f_{p}(\sigma)$. To remove this indeterminacy and enhance optimality, extra planning objectives are introduced while the optimal path cost is imposed as a constraint. In particular, for (17), the optimal path is found in a length-regulated form as
[…]
@@ 4.2 Global passage width PTOPP
Despite its simplicity and versatility, only taking into account $f_{p}(\sigma)$ may not reach high-quality paths for ignoring passages other than $f_{p}(\sigma)$. For example, $f_{p}(\sigma)$ is an inevitable passage from the start $\mathbf{x}_{0}$ in Figure 8. MPW-PTOPP essentially degrades to shortest path planning. This further scales to occasions where feasible paths have inevitable passages. A typical situation is enclosure of $\mathbf{x}_{0}$ by passages. Paths must traverse one surrounding passage (see Figure 8). Outer passages wider than them are muted in costs, highlighting the limitation of MPW-PTOPP. To alleviate this degradation, a comprehensive coverage of $P_{\sigma}$ is required. Conceptually, if traversed passage widths more than $f_{p}(\sigma)$ in $P_{\sigma}(1)$ are incorporated in path costs, problems are called global passage width PTOPP (GPW-PTOPP).
Cost formalisms in GPW-PTOPP are non-trivial to involve multiple passages properly. Some common measures are not directly applicable as costs. For instance, maximizing the sum of passed passage widths incurs an ergodic passage traversal, and the average passage width lacks monotonicity. For better comparability of passage widths, denote $\mathbf{p}_{\sigma}(\tau)$ the passed passage widths in ascending order and $\bar{P}_{\sigma}(\tau)$ the sorted permutation of $P_{\sigma}(\tau)$ such that $\|\bar{P}_{\sigma}(\tau,i)\|\leq\|\bar{P}_{\sigma}(\tau,j)\|$ for $i<j$, then
@@
where $\mathbf{p}_{\sigma}(\tau,i)=\|\bar{P}_{\sigma}(\tau,i)\|$. To involve all these traversed passage widths in costs, the path cost summing them up is framed as
@@
where $\mathbf{w}_{p}\succeq\mathbf{0}$ with nonnegative entries is the weight vector. How $f_{c}(\sigma)$ updates is the key to compatibility, but it is not well-defined since the dimension of $\mathbf{p}_{\sigma}$ generally increases in path concatenation. An important step is to adopt a fixed and large dimension of $\mathbf{p}_{\sigma}$. Say the dimension is q, $\mathbf{p}_{\sigma}$ is initialized as $\bar{\varepsilon}_{p}\mathbf{1}_{q}$ where $\bar{\varepsilon}_{p}$ is a large value as in (15). $\mathbf{1}_{q}$ is the q-dimensional all-one vector. Every sorted passage width is now assigned a weight. Elements in $\mathbf{w}_{p}$ are in descending order so that narrower passages have higher priorities to be avoided, i.e., $\mathbf{w}_{p}=[w_{p,1},w_{p,2},...,w_{p,q}]^{\mathsf{T}}$, $w_{p,i}<w_{p,i+1}$.
Figure 8: On the left, the start and target are on two sides of the narrowest passage. Feasible paths must traverse this passage. The start position on the right is enclosed by passages. Any feasible path will pass one of them.
While $\mathbf{p}_{\sigma}$ now preserves its dimensionality and (21) is monotonically non-increasing, it is easy to verify that $f_{c}(\cdot)$ is not order-preserving as a sum of paired $w_{p,i}\mathbf{p}_{\sigma}(1,i)$. To resolve this, the lexicographic order is utilized for comparing $\mathbf{p}_{\sigma}$ and $f_{c}(\cdot)$ is specially designed to preserve this order via configuring $\mathbf{w}_{p}$. Formally, if $\mathbf{p}_{\sigma_{1}}$ is lexicographically smaller than $\mathbf{p}_{\sigma_{2}}$, then
@@
When $\mathbf{p}_{\sigma_{1}}\prec_{\text{lex}}\mathbf{p}_{\sigma_{2}}$, the narrowest passage $\sigma_{2}$ passes is wider than $\sigma_{1}$ except those common passages for $\sigma_{1},\sigma_{2}$. A smaller cost is thus expected from $f_{c}(\sigma_{2})$. This can be done by setting large differences between successive $\mathbf{w}_{p}$ items so that $w_{p,i}/w_{p,i+1}\gg 1$. For a given passage distribution, by setting sufficiently large differences between $w_{p,i}$, the following holds
@@
In practice, q can be fairly small to consider top k narrowest passages and such problems are called k-minimum passage width PTOPP (k-MPW-PTOPP), as a generation of (17). Directly composing a compatible trade-off cost analogous to (14) is attainable as $f_{c}(\sigma)=\text{Len}(\sigma)-\mathbf{w}^{\mathsf{T}}\mathbf{p}_{\sigma}(1)$. However, weights in $\mathbf{w}$ are set equal to be order-preserving, rendering it unable to prioritize avoiding narrower passages (see Appendix A.2 for more analysis of compatible GPW-PTOPP path cost design).